can you help me solve: x+2y=22
-x+2y=10 and the answer is with an ordered pair!

Thank you!

Answer by stanbon(75887)

solve:
x+2y=22
-x+2y=10
------
Add and solve for "y":
4y = 32
y = 8
------
Solve for "x":
x + 2y = 22
x + 16 = 22
x = 6
----
Ans: (6,8)
=======
Cheers,
Stan H.
============